Grade 12th PassAnalytical Geometry

If ax2+2bxy+bx2+2x+2y+c=0 represents a circle, then the true statement among the following is
  1. a+b+c > 6
  2. a+b+c > 4
  3. a+b+c 3
  4. a+b+c > 5
pls answer with a reason............
& pls answer fastly.................

Charitha Gorrela
9 Years agoGrade 12th Pass

1 Answer

Vikas TU
9 Years ago
Dear Student,
The two lines given are r = (6i + 2j + 2k) + t(i – 2j + 2k) and r = (- 4i – k) + s(3i – 2j – 2k).
For two lines x = a + tb and y = c + sd, the shortest distance between them are 
d = | bxd/|bxd| . (c – a) | = |(8i + 8j + 4k).(-10i – 2i – 3k)/12| = |(- 80 – 16 – 12)/12| = 108/12 = 9.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)